# Self-Critique and Refinement for Faithful Natural Language Explanations

### v1

@@ 4.5 Experimental Setups
We use greedy decoding throughout our pipeline and experiment with up to 3 refinement rounds. For attribution-based IWF, we compare two attribution methods: (1) gradient-based attribution using Integrated Gradients (IWF-IG; Sundararajan et al. (2017)), identified as the most faithful post-hoc explanations Atanasova et al. (2020a), and (2) attention-based attribution (IWF-Attn) leveraging the model’s attention mechanisms. A more detailed description of these attribution methods is provided in Appendix D. For all important-word feedback variants, we use the top-5 important words as feedback. For the SC-NLE baseline, we sample 20 candidate explanations with temperature 1.0 and select the most representative one using semantic centroid voting, as described in Section 4.4. Detailed ablation studies on various parameters are provided in Appendix E.
@@
Our entire pipeline operates in a zero-shot setting, with stage-specific instructions designed for each dataset. Complete prompt templates are provided in Appendix H.
5 Results
5.1 Main Results
@@
Table 1 presents our comprehensive evaluation results after 3 refinement rounds. Additional results from intermediate refinement rounds, complementary metrics (prediction accuracy and counter rates), and detailed visualizations are provided in Appendix F.
@@
The SR-NLE framework shows superior performance over baseline methods in most experimental settings. Our best implementation (IWF-Attn) reduces unfaithfulness rates by an average of 18.79% compared to Init-NLE and 13.72% compared to SC-NLE. Even our least effective method (NLF), despite underperforming in isolated cases (e.g., e-SNLI with Falcon), still achieves an average reduction of 7.23% compared to Init-NLE and 2.16% compared to the SC-NLE baseline, demonstrating the overall effectiveness of our framework.
@@
All three IWF implementations consistently outperform NLF across all experimental settings. On average, IWF-Attn, IWF-IG, and IWF-Pmt, achieve 11.56%, 11.28%, and 10.35% lower unfaithfulness rates than NLF, respectively. This performance gap demonstrates that explicit important word feedback provides more effective guidance for refinement than natural language feedback.
Prompt-based IWF comparable to attribution-based IWF
A notable finding is that prompt-based IWF performs nearly identically to attribution-based implementations, with average unfaithfulness rates of 36.02% (IWF-Attn), 36.30% (IWF-IG), and 37.23% (IWF-Pmt), differing by only 1.21%. This suggests that language models possess inherent capabilities to identify important reasoning components through carefully designed prompts without requiring specialized attribution techniques. Such capability has significant practical value, as prompt-based methods are considerably more efficient and accessible than traditional attribution methods.
@@
While the results in Table 2 suggest that SR-NLE substantially outperforms the state-of-the-art G-TEX in terms of explanation faithfulness, this comparison is not fully controlled. The reported numbers correspond to each method’s best-performing configuration on each dataset, and differences in counterfactual generation strategies or data splits may influence the outcomes. Nevertheless, the results provide a useful reference point, demonstrating the potential of SR-NLE as a lightweight and effective alternative for improving explanation faithfulness, compared to G-TEX that requires architectural changes and additional fine-tuning.
@@ 5.3 Detailed Analysis
In this section, we conduct in-depth analyses to better understand the effectiveness of our SR-NLE framework, examining faithfulness state transitions, refinement efficiency across rounds, and the relationship between explanation length and faithfulness.
@@
Figure 5: Unfaithfulness rates across successive refinement rounds for different feedback methods, averaged across 12 model-dataset combinations. The steepest decline occurs in the first round, with progressively smaller reductions in subsequent rounds.
@@
Figure 6: Unfaithfulness rates versus explanation lengths across refinement rounds for different feedback methods, averaged across 12 model-dataset combinations.
@@ 6 Conclusion
In this work, we presented SR-NLE, a framework for improving the faithfulness of NLEs through an iterative self-critique and refinement process. By enabling LLMs to iteratively refine their own explanations with appropriate self-feedback, our approach significantly reduces unfaithfulness rates across multiple datasets and models without requiring external supervision, additional training or architectural changes. Our experiments demonstrate that important-word feedback consistently outperforms natural language feedback, with attention-based methods achieving the best results. The detailed analyses reveal that our framework efficiently targets critical reasoning components, successfully converts unfaithful explanations to faithful ones, and optimizes explanation content rather than merely increasing length. These findings suggest that self-refinement offers a promising path toward more faithful and trustworthy explanation generation. Future work could explore additional feedback mechanisms and investigate the applicability of SR-NLE to broader domains and more diverse reasoning tasks.
@@
We rely on counterfactual tests as the sole evaluation method for measuring explanation faithfulness. While this metric offers objective signals aligned with our goal, it reflects only one type of faithfulness criterion. Future work could explore additional automatic tests—such as consistency tests, simulatability tests—to provide a more comprehensive view of explanation faithfulness. Moreover, our evaluation approach does not capture other important aspects of explanation quality such as plausibility, completeness, or alignment with human-annotated references.
@@ Appendix A Datasets
For our experiments, we selected the first 1,000 instances from each dataset’s test set. The full test set of ComVE contains 1,000 instances, while ECQA and e-SNLI have 2,194 and 9,824 instances respectively. To verify the representativeness of these subsets, we analyzed their label distributions compared to the full test sets, as shown in Table 3. The slight deviation from a 50/50 split in ComVE results from the random ordering process during dataset preparation, where we randomly positioned the sentence that violates common sense as either the first or second sentence.
@@ Appendix E Ablation Studies
All ablation studies are conducted on the same 100-instance subsets sampled from each dataset. For the IG integration steps (Section E.1), we directly use these 100 original instances. For the other two ablation studies (Section E.2 and E.3), we follow the same procedure as in the main experiments to compute unfaithfulness: for each of the 100 instances, we generate 20 interventions (totaling 2000 intervened instances), and then select and perform experiments on the counter instances. It’s important to note that we conducted these ablation studies across all dataset and model combinations. Therefore, all metrics reported in these sections represent averaged values across the entire experimental matrix, providing a comprehensive view of our method’s performance across different conditions.
@@ E.1 Integration Steps for IG Attribution
We investigated the impact of integration steps on the convergence of IG attribution calculations by experimenting with nine different step settings ranging from 100 to 1500.
Figure 7 shows that most models (Falcon, Llama, and Mistral) reach reasonable convergence around 500 integration steps, with the convergence delta showing minimal changes beyond this point. In contrast, the Qwen model exhibits higher variability and slower convergence, requiring more steps to achieve stable attribution values. This may be caused by architectural differences that affect how gradients are calculated and propagated through the model. Based on these observations, we selected 500 integration steps for Falcon, Llama, and Mistral, while using 1000 steps for Qwen in our main experiments.
@@ E.2 SC-NLE Parameters
We investigated the impact of two key parameters for our SC-NLE baseline: candidate explanation count and sampling temperature.
Figure 8 shows that increasing the number of candidates reduces unfaithfulness, with significant improvements up to 20 samples. Temperature also affects performance, with temperature 1.0 consistently outperforming lower values, especially at higher sample counts. Based on these results, we selected 20 candidate explanations at temperature 1.0, balancing performance improvements with computational cost.
E.3 Number of Important Words
@@
We investigated the optimal number of important words (Top-N) for our Important-Word Feedback through two complementary analyses: examining unfaithfulness changes and analyzing attribution distribution patterns.
Figure (9a) shows unfaithfulness decreases as N increases from 1 to 9 across all feedback types, with the most significant improvements occurring between N=1 and N=5. Concurrently, Figure (9b) reveals that the top-5 important words capture approximately 70-80% of the total attribution magnitude, despite typically representing only a small fraction of input tokens. Therefore, we selected N=5 for all our main experiments.
@@ F.1 Prediction Accuracy
Dataset Falcon Llama Mistral Qwen
@@ F.2 Counter Rates
Dataset Falcon Llama Mistral Qwen
@@
F.3 Intermediate Results
@@
Table 8: Unfaithfulness rates (%) after refinement rounds 1 (R1) and refinement round 2 (R2).
F.4 Performance Visualization
@@ Appendix G Algorithms
Algorithms 1-3 present the SR-NLE framework, attribution-based important-word feedback, and semantic centroid voting strategy for SC-NLE, respectively.
@@
2:Final answer $y$ and refined explanation $e^{k}$
3:$y\leftarrow\mathcal{M}(p_{\text{ans}}\oplus x)$
4:$e^{0}\leftarrow\mathcal{M}(p_{\text{exp}}\oplus x\oplus y)$
@@
9:         $\mathcal{S}\leftarrow\textsc{AttributionScore}(\mathcal{M},p_{\text{ans}},x,y)$triangleright Algorithm 2
@@
Algorithm 2 Attribution-based IWF
@@
4:for each token $y_{j}$ in answer span do $\triangleright$ Sequential token attribution
5:     for each token $x_{i}$ in full model input do $\triangleright$ Token-level attribution
@@
8:     $a_{i}\leftarrow\sum_{j=1}^{|y|}a_{i,j}$triangleright Target-level aggregation
@@
10:     $\text{S}(w)\leftarrow\sum_{i\in\text{indices}(w)}a_{i}$triangleright Word-level aggregation
@@
3:$embeddings\leftarrow M.encode(E)$triangleright Encode all explanations
4:$centroid\leftarrow\frac{1}{n}\sum_{i=1}^{n}embeddings_{i}$triangleright Compute centroid embedding
5:$similarities\leftarrow[\,]$
@@
7:     $similarities_{i}\leftarrow\text{cosine\_similarity}(centroid,embeddings_{i})$triangleright Compute similarity to centroid
8:$best\_idx\leftarrow\arg\max(similarities)$triangleright Select most similar to centroid

### v2

@@ 4.5 Experimental Setups
We use greedy decoding throughout our pipeline and experiment with up to $K=3$ refinement rounds. For attribution-based IWF, we compare two attribution methods: (1) gradient-based attribution using Integrated Gradients (IWF-IG; Sundararajan et al. (2017)), identified as the most faithful post-hoc explanations Atanasova et al. (2020a), and (2) attention-based attribution (IWF-Attn) leveraging the model’s attention mechanisms. A more detailed description of these attribution methods is provided in Appendix D. For all important word feedback variants, we use the top-5 important words as feedback. For the SC-NLE baseline, we sample 20 candidate explanations with temperature 1.0 and select the most representative one using semantic centroid voting, as described in Section 4.4. Detailed ablation studies on various parameters are provided in Appendix E.
@@
Our entire pipeline operates in a zero-shot setting, with stage-specific instructions designed for each dataset. Complete prompt templates are provided in Appendix I.
@@
5 Results
5.1 Main Results
Table 1 presents our comprehensive evaluation results after 3 refinement rounds. Additional results from intermediate refinement rounds, complementary metrics, additional analysis, and detailed visualizations are provided in Appendix F.
@@ 5.1 Main Results
The SR-NLE framework shows superior performance over baseline methods in most experimental settings. Our best implementation (IWF-Attn) reduces unfaithfulness rates by an average of 18.79% compared to Init-NLE and by an average of 13.72% compared to SC-NLE. Even our least effective method (NLF), despite underperforming in isolated cases (e.g., e-SNLI with Falcon), still achieves an average reduction of 7.23% compared to Init-NLE and 2.16% compared to the SC-NLE baseline, demonstrating the overall effectiveness of our framework.
@@
All three IWF implementations consistently outperform NLF across all experimental settings. On average, IWF-Attn, IWF-IG, and IWF-Pmt achieve 11.56%, 11.28%, and 10.35% lower unfaithfulness rates than NLF, respectively. This performance gap demonstrates that explicit important word feedback provides more effective guidance for refinement than natural language feedback.
Comparable performance across IWF variants
A notable finding is that prompt-based IWF performs similarly to attribution-based implementations, with average unfaithfulness rates of 36.02% (IWF-Attn), 36.30% (IWF-IG), and 37.23% (IWF-Pmt), differing by only 1.21%. This suggests that the IWF framework is robust to different word selection strategies, with prompt-based methods offering practical advantages in terms of efficiency, accessibility, and reliability (3.75% hallucination rate; see Appendix F.5). To better understand this robustness, we also conducted additional experiments exploring the impact of word selection quality on IWF performance (detailed in Appendix F.6).
@@
While the results in Table 2 suggest that SR-NLE substantially outperforms the state-of-the-art G-TEX in terms of explanation faithfulness, this comparison is not fully controlled. The reported numbers correspond to each method’s best-performing configuration on each dataset, and differences in counterfactual generation strategies or data splits may influence the outcomes. Nevertheless, the results provide a useful reference point, demonstrating the potential of SR-NLE as a lightweight and effective alternative for improving explanation faithfulness, compared to G-TEX, which requires architectural changes and additional fine-tuning.
@@ 5.3 Detailed Analysis
In this section, we conduct further in-depth analyses to better understand the effectiveness of our SR-NLE framework.
@@
Figure 5: Unfaithfulness rates across successive refinement rounds for feedback methods, averaged across 12 model-dataset combinations.
@@
Figure 6: Unfaithfulness rates versus explanation lengths across refinement rounds for feedback methods, averaged across 12 model-dataset combinations.
@@
Original Question: There was only one room in the place where Bill slept. It had a bed, a fridge, a stove, a couch, and a television. Where might he be?
Answer options: (A) motel (B) school (C) hotel (D) apartment (E) friend’s house
Model prediction: (D) apartment
Intervened Question: There was only one cozy room in the place where Bill slept. It had a bed, a fridge, a stove, a couch, and a television. Where might he be?
Answer options: (A) motel (B) school (C) hotel (D) apartment (E) friend’s house
Model prediction: (A) motel
Initial Explanation: Bill’s room contains amenities typical of a motel, such as a bed, fridge, stove, couch, and television, which are not usually found together in a school setting. Unfaithful
Natural Language Feedback (NLF) Important-Word Feedback (IWF-Attn)
Feedback 1: The explanation accurately reflects the reasoning behind selecting option (A) as it highlights the motel-like amenities present in Bill’s room. However, it could be improved by explicitly mentioning that the combination of a bed, fridge, stove, couch, and television is more characteristic of a motel or hotel room than other options like a school or apartment. This would make the reasoning clearer and more convincing. No other significant improvements are needed. Feedback 1: The 5 most important words that contributed to your prediction are: one, a, cozy, be, there.
Refinement 1: Bill’s room contains a bed, fridge, stove, couch, and television, which are typical amenities found in a motel or hotel room, not in a school or apartment setting. This combination of features makes option (A) Motel the most appropriate choice. Unfaithful Refinement 1: Bill’s room, described as cozy with amenities like a bed, fridge, stove, couch, and television, suggests a motel setting, as these features are commonly found together in such accommodations. Faithful
Refinement after 3 rounds: Bill’s room, equipped with a bed, fridge, stove, couch, and television, is indicative of a motel or hotel room, where such amenities are common for temporary accommodation and comfort. The presence of a bed, fridge, and stove suggests a living space, which aligns more with a motel setting than a school or apartment. This combination of features makes option (A) Motel the most appropriate choice. Unfaithful Refinement after 3 rounds: Bill’s cozy room, complete with a bed, fridge, stove, couch, and television, suggests a motel setting, where such a combination of amenities is commonly found. Faithful
Table 3: Case study comparing NLF and IWF-Attn (our best-performing variant) on the ECQA dataset. The intervened word, highlighted in blue, successfully changes the model prediction. Faithful indicates the explanation/refinement is faithful as judged by the counterfactual test, while Unfaithful indicates the opposite.
Case Study
Table 3 presents a case study from the ECQA dataset comparing NLF and IWF-Attn. Starting from the same unfaithful initial explanation, IWF-Attn successfully achieves faithfulness after one refinement round guided by its identification of the five most important words for the prediction. In contrast, NLF fails to achieve faithfulness even after three rounds of refinement. Despite receiving detailed feedback suggesting various improvements, NLF’s refinements become progressively longer but still remain unfaithful. Complete refinement details and additional examples from e-SNLI and ComVE can be found in Appendix H.
@@ 6 Conclusion
In this work, we presented SR-NLE, a framework for improving the faithfulness of NLEs through an iterative self-critique and refinement process. By enabling LLMs to iteratively refine their own explanations with self-feedback, our approach significantly reduces unfaithfulness rates across multiple datasets and models without requiring external supervision, additional training or architectural changes. Our experiments demonstrate that IWF consistently outperforms NLF, with attention-based methods achieving the best results. The detailed analysis reveals that our framework efficiently targets critical reasoning components, successfully converts unfaithful explanations to faithful ones, and optimizes explanation content rather than merely increasing length. These findings suggest that self-refinement offers a promising path toward more faithful explanation generation. Future work could explore additional feedback mechanisms and investigate the applicability of SR-NLE to a broader set of domains and diverse reasoning tasks.
@@
We rely on counterfactual tests as the sole evaluation method for measuring explanation faithfulness. While this metric offers objective signals aligned with our goal, it reflects only one type of faithfulness criterion. Future work could explore additional automatic tests–such as consistency tests and simulatability tests–to provide a more comprehensive view of explanation faithfulness. Moreover, our evaluation approach does not capture other important aspects of explanation quality, such as plausibility, completeness, or alignment with human-annotated references.
@@ Appendix A Datasets
For our experiments, we selected the first 1,000 instances from each dataset’s test set. The full test set of ComVE contains 1,000 instances, while ECQA and e-SNLI have 2,194 and 9,824 instances, respectively. To verify the representativeness of these subsets, we analyzed their label distributions compared to the full test sets, as shown in Table 4. The slight deviation from a 50/50 split in ComVE results from the random ordering process during dataset preparation, where we randomly positioned the sentence that violates common sense as either the first or second sentence.
@@ Appendix E Ablation Studies
All ablation studies are conducted on the same 100-instance subsets sampled from each dataset. For the IG integration steps (Section E.1), we directly use these 100 original instances. For the other two ablation studies (Section E.2 and E.3), we follow the same procedure as in the main experiments to compute unfaithfulness: for each of the 100 instances, we generate 20 interventions (totalling 2000 intervened instances), and then select and perform experiments on the counter instances. It’s important to note that we conducted these ablation studies across all dataset and model combinations. Therefore, all metrics reported in these sections represent averaged values across the entire experimental matrix, providing a comprehensive view of our method’s performance across different conditions.
@@ E.1 Integration Steps for IG Attribution
We investigated the impact of integration steps on the convergence of IG attribution calculations by experimenting with nine different step settings ranging from 100 to 1500. Figure 7 shows that most models (Falcon, Llama, and Mistral) reach reasonable convergence around 500 integration steps, with the convergence delta showing minimal changes beyond this point. In contrast, the Qwen model exhibits higher variability and slower convergence, requiring more steps to achieve stable attribution values. This may be caused by architectural differences that affect how gradients are calculated and propagated through the model. Based on these observations, we selected 500 integration steps for Falcon, Llama, and Mistral, while using 1000 steps for Qwen in our main experiments.
@@ E.2 SC-NLE Parameters
We investigated the impact of two key parameters for our SC-NLE baseline: candidate explanation count and sampling temperature. Figure 8 shows that increasing the number of candidates reduces unfaithfulness, with significant improvements up to 20 samples. Temperature also affects performance, with temperature 1.0 consistently outperforming lower values, especially at higher sample counts. Based on these results, we selected 20 candidate explanations at temperature 1.0, balancing performance improvements with computational cost.
E.3 Number of Important Words
@@
We investigated the optimal number of important words (top-$N$) for our Important Word Feedback through two complementary analyses: examining unfaithfulness changes and analyzing attribution distribution patterns. Figure (9a) shows that unfaithfulness decreases as $N$ increases from 1 to 9 across all feedback types, with the most significant improvements occurring between $N=1$ and $N=5$. Concurrently, Figure (9b) reveals that the top-5 important words capture approximately 70-80% of the total attribution magnitude, despite typically representing only a small fraction of input tokens. Therefore, we selected $N=5$ for all our main experiments.
@@ F.1 Prediction Accuracy
Falcon Llama Mistral Qwen
@@ F.2 Counter Rates
Falcon Llama Mistral Qwen
@@
F.3 Average Sequence Lengths
Table 9 reports the average sequence lengths of counter instances for each model-dataset combination. We provide results under two settings: Full, which denotes the total word length of the input, and Unique, which denotes the word length after removing duplicate words. The lengths range from about 13 to 21 words under the Full setting, while the Unique setting is consistently shorter, around 9 to 16 words.
Falcon Llama Mistral Qwen
Full
ComVE 15.72 15.13 14.96 16.18
ECQA 12.99 13.53 13.45 13.64
e-SNLI 21.08 21.55 20.70 21.35
Unique
ComVE 10.16 9.35 9.67 10.26
ECQA 12.25 12.66 12.55 12.82
e-SNLI 15.65 16.28 15.78 15.81
Table 9: Average sequence lengths of counter instances. Full refers to the total word length, while Unique refers to the word length after removing duplicate words.
F.4 Intermediate Results
@@
Table 10: Unfaithfulness rates (%) after refinement rounds 1 (R1) and refinement round 2 (R2). Best (lowest) results per dataset-model combination are bolded, second best are underlined.
F.5 Reliability of Prompt-based Important Words Selection Strategy
To assess the reliability of prompt-based important words selection, we examined whether the important words identified by IWF-Pmt are grounded in the input. Specifically, we measured the hallucination rate, defined as the proportion of top-5 selected words that do not appear in the input. Table 11 presents the results across all dataset-model combinations. While we observe some variation—with ECQA showing slightly higher rates and Qwen exhibiting more hallucination compared to other models—the overall average hallucination rate across the 12 combinations is only 3.75%. This low rate demonstrates that prompt-based word selection is highly reliable, with the selected words being well-grounded in the input.
Falcon Llama Mistral Qwen
ComVE 0.01 0.01 0.01 0.05
ECQA 0.07 0.10 0.03 0.08
e-SNLI 0.01 0.01 0.02 0.05
Table 11: Hallucination rates (%) of extracted important words in IWF-Pmt, measured as the proportion of top-5 words not appearing in the input.
F.6 Word Selection Quality Analysis
To assess the impact of word selection quality on IWF performance, we conducted experiments comparing different selection methods against a random baseline. Specifically, we randomly selected five words from the input and applied three rounds of refinement on the e-SNLI dataset across all four models, averaging results over three random seeds. As shown in Table 12, while the random baseline shows slightly higher unfaithfulness rates, it achieves performance close to both prompt-based and attribution-based IWF methods. To understand this result, we analyzed how often the intervened word (i.e., the true reasoning factor that made the label change) appeared in the top-$N$ selected words. Table 13 shows that current selection methods—whether prompt-based or attribution-based—capture the true reasoning word at rates similar to random selection. These findings reveal two important insights:
  1. 1.
    Robustness of IWF: Even with suboptimal word selection, IWF can effectively improve explanation faithfulness, demonstrating that the iterative refinement process in SR-NLE contributes significantly to performance improvements beyond the quality of word attribution.
  2. 2.
    Opportunities for improvement: Current word attribution methods have considerable room for enhancement. As better attribution techniques are developed, IWF could potentially achieve even stronger performance within our SR-NLE framework.
This analysis underscores that IWF’s effectiveness stems from the iterative refinement process in our SR-NLE framework rather than perfect word identification, making IWF a robust feedback mechanism that can benefit from future advances in attribution methods.
Falcon Llama Mistral Qwen Avg.
Random
21.18
  $\pm$0.26
35.32
  $\pm$0.71
37.68
  $\pm$0.63
22.91
  $\pm$0.68
29.28
  $\pm$7.64
IWF-Pmt 22.01 36.16 37.80 24.43 30.10
IWF-Attn 18.21 34.94 38.49 19.10 27.69
IWF-IG 18.35 37.38 35.86 21.98 28.39
Table 12: Comparison of random baseline against IWF methods on e-SNLI dataset (unfaithfulness rates in %). Random results show mean ± standard deviation over three seeds. All methods use three refinement rounds.
Top-1 Top-2 Top-3 Top-4 Top-5
Random 6.58 13.28 20.12 27.03 34.00
Pmt 3.25 8.70 14.89 22.17 30.51
Attn 4.52 14.84 25.52 36.07 46.25
IG 8.58 17.14 25.22 32.70 39.87
Table 13: Intervened word inclusion rate (%) in the top-$N$ selected words under the random baseline and IWF methods.
F.7 Performance Visualization
@@ Appendix G Algorithms
Algorithms 1-3 present the SR-NLE framework, attribution-based IWF Score, and semantic centroid voting strategy for SC-NLE, respectively.
@@
1:Model $\mathcal{M}$, prompts $\{p_{\text{ans}},p_{\text{exp}},p_{\text{fb}},p_{\text{ref}}\}$, input $x$, feedback type $t\in\{\text{NLF},\text{IWF}\}$, IWF method $m\in\{\text{prompt-based},\text{attribution-based}\}$, refinement rounds $K$, number of important words NN
2:Final answer $y$ and refined explanation $e^{K}$
@@
[…]
@@
9:   $\mathcal{S}\leftarrow\textsc{AttributionScore}(\mathcal{M},p_{\text{ans}},x,y)$ $\triangleright$ Algorithm 2
@@
Algorithm 2 Attribution-based IWF Score
@@
4:for each token $y_{j}$ in answer span do $\triangleright$ Sequential target token attribution
5:  for each token $x_{i}$ in full model input do $\triangleright$ Token-level computation
@@
8:  $a_{i}\leftarrow\sum_{j=1}^{|y|}a_{i,j}$ $\triangleright$ Target-level aggregation
@@
10:  $\text{S}(w)\leftarrow\sum_{i\in\text{indices}(w)}a_{i}$ $\triangleright$ Word-level aggregation
@@
[…]
@@
7:  $similarities_{i}\leftarrow\text{cosine\_similarity}(centroid,embeddings_{i})$ $\triangleright$ Compute similarity to centroid
[…]